Meeting notes:

Eric -- Working hard on artdaq-otsdaq interface overhaul. Working on error handling.

John -- Fiber cassettes and patch panels should be arriving this week. Fireflies in a couple weeks. Ordered standoffs/hardware for Timing cards. We have a spreadsheet for Timing Cards - should add columns for Firefly serial numbers. Tag DTCs and Timing cards separately for PREP. STM will use “early production” - likely have 3 spares.

Antonio -- Been using mu2eshift user environment and getting up to speed on DCS/EPICS. Currently noticing values are not being refreshed or we are not forcing values properly. Working on this. Dan’s last day is today.

Micol -- Fixed ROC core firmware retransmission. Understood first retransmission request - both DTC and ROC should be reset. Almost everything seems to work. Stress test with a few iterations works. Looking at loopback with Greg.

Greg -- Configuring CFO also now, seems to work. With DTC and ROC. Now want to send loopback marker from CFO -> DTC -> ROC and back.

Rick -- Fixed a few CRV related bugs - ready to test with them. Started working on event building testing.

Ron -- Full file path comes in next artdaq release.

Ryan -- Thinking about last timing card.

Vivian -- Simulation file running through artdaq board reader to ROOT file. Then tried to connect hardware and ran against the new release waiting. Trying to use a chant to do a data request.

Bertrand -- Reco models changing.. May cut 1ms from processing. Currently at “3.4ms”

Sophie -- Waiting for STM DTC update.

Stefan - With Oliver. Will get up to speed on something next week. Oliver will be here for 3 more weeks. Work with Antonio.